These are the news starting 01 Jan 2007
Click here for the news of the previous years.
|25 Dec 07||CO2 sensor definitively out of order!|
|23 Dec 07||The membrane pump of the CO2 sensor broke down. As I have no spares, a different pump from an old disused sensor has been installed as a kludge. Seems working! The last week CO2 readings should be considered suspect!|
|22 Dec 07||Recalibration and maintenance of CO2 sensor (09:00)|
|21 Dec 07||Recalibration and maintenance of NOx and O3 sensors (10:00)|
|07 Dec 07||
1. Uploaded Nov07 data to archive.
2. The almost 540 ppm CO2 spike of 03Dec07 is an artefact: the sampling tube went loose and I spent an hour in the sensor room, so this spike is a fingerprint of my respiration. Correct spike is probable 450 ppm. Will be corrected in the monthly data file when available.
|05 Dec 07||Our traditional Christmas logo is back. Thanks to Colette Heirendt for this artistic work.|
|03 Dec 07||Attention: I forgot to switch back the NO/NO2 sensor to the ug/m3 scale after calibration (which is done in ppb). Correct all readings from 22Nov07 07:30 to 03Dec07 14:28 by multiplying the NO data by 1.34 and the NO2 by 2.05. Readings are again ok starting 03Dec07 15:00 UTC. The monthly data file nov07.dat will contain this message..|
|22 Nov 07||07:45 UTC: Recalibration of NOx sensor. Rechecked and recalibrated CO2 sensor|
|21 Nov 07||Maintenance and recalibration work on all gas sensors from 14:10 on|
|10 Nov 07||Restarted new all.dat file with the raw logger data (begin: 10Nov07 00:00). The previous file is called all_01Jul06to09Nov.dat., located in the data archive.|
|09 Nov 07||Uploaded Oct07 data to data archive.|
|31 Oct 07||Repaired pump broke down again on 29th October, due to damaged bearing. As a new pump would cost well over 1000 Euro, we did a definitive repair in-house (thanks to technician Jean-Paul Klein for helping out) by replacing the worn-out SKF bearing by a new one, at a cost of 5 Euro ! So even if meteoLCD saved a lot of money on this, it does not mean that we are rich!|
|22 Oct 07||
1. Main pump of ground ozone sensor breakdown the
21th Oct. 07. I will try to replace the pump by an older
2. General main power loss for the whole Diekirch region at approx. 09:00, duration less than 30 minutes
|16 Oct 07||Recalibrated CO2 sensor.|
|14 Oct 07||WBGT is not measured anymore during the cold season. Page with live-data adapted to show this. (WBGT measurements should be discarded from 01Oct07 on)|
|12 Oct 07||Maintenance work by Envitec. NOx sensor repaired and back to work. All gas sensors recalibrated.|
|07 Oct 07||Uploaded Sep07 data into data archive.|
|27 Sep 07||Uploaded Aug07 data into data archive. Sorry for the delay, was very busy with our exhibition/workshop of vintage mechanical calculators (look at http://computarium.lcd.lu). Remember that not yet archived data are usually available through the "all data since logger restart" link.|
|13 Sep 07||The Dobson data have been updated to include the measurements made by Marcel Thilmany from 25 july to 30 Aug. during my leave. Instrument was Microtops #3012, and location was Insenborn (lat. 49,32°, long. 5.32°E, elev. 367m). The raw readings have been divided by 0.95 to make them comparable to those made by the usual #5375 instrument.|
|18 Aug 07||
1. NOx sensor bad since several weeks. Shut off
today. Repair not possible before 2 months
2. There is also an internal malfunction of the O3 sensor (an internal temperature is too low). This has possibly no impact on the measurements.
3. Recalibration done today morning
4. Upload of jun07.dat and jul07.dat to data archive
|02 Jul 07||Uploaded all_01Jul06to30Jun07.dat file into data archive. This huge file contains all the raw untouched logger data for these 12 months. See the "ASCII DATA TODAY" link from the "Today's Meteorological Data" page to find the column definitions.|
|15 Jun 07||
Recalibration of gas sensors from 13:00 to 15:00
UTC. CO2 calibrated with new gas bootle, containing CO2 at 398.3
(+/-1 1%) what is the (expensive!) Primary Standard of Praxair.
*** Attention: From 08Jun to 15Jun units of NO/NO2 were ppb instead of ug/m3. To revert to ug/m3. multiply NO2 readings by 1.34 and NO2 readings by 2.05. Correct ug/m3 units restated 15Jun07 15:00 ***
|10 Jun 07||Uploaded may07.dat file to data archive.|
|08 Jun 07||Maintenance work by Envitec on O3, CO2 and NOx sensors. No recalibration.|
|25 May 07||A severe thunderstorm with golf ball sized hail passed over Diekirch and surrounding region during 20.00 to 22:00 UTC. Maximum activity was between 21:00 and 22:00. Watch movies from 20070525 and 20070526, accessible from the live data page. By good luck, all our sensors survived unharmed!|
|24 May 07||
1. WBGT sensor ran dry yesterday (23/05/07);
distilled water reservoir for wetbulb refilled today at09:00 UTC.
2. Maintenance and recalibration work
|17 May 07||
Added a link to the exported data from our Davis
Vantage Pro backup station on the main page with the live
data. The text files are stored as ddmmyy.txt and typically hold
2 days of data. Be aware: the internal date format in these
files is "american style" mmddyy ! No pictures nor
graphs, only bare data!
The readme.html file found under "Equipment" has been updated to version 5.02
|16 May 07||Bird droppings have clogged the rain-sensor, so we missed the last days. Rainfall data from our backup station will be used in the monthly data file.|
Webcam functional again.
April data uploaded (apr07.dat)
|30 Apr 07||
From 12:00 on: maintenance work and
Attention: webcam is out of order, due to a software problem!
|24 Apr 07||CO2 sensor reconnected at 06:55 to usual sampling line (about 25m above ground).|
|21 Apr 07||CO2 is still being sampled from about 3m height above ground; the lesser diurnal amplitude could be a result of this, but it could also be caused by possible problems with the pumping system.|
|16 Apr 07||Started at 13:00 to sample CO2 from a tube at a height of about 3m above ground level. The system is very experimental; the experiment is planned to run for 2 weeks, but might be aborted in case of problems..|
|14 Apr 07||Restarted WBTG (heatstress) sensors at 16:30|
|05 Apr 07||Uploaded mar07.dat file into data archive.|
|25 Mar 07||There was a computer glitch sometimes after the 21th March. Ok again, no data lost. Strange lines in the plots should be discarded.|
|17 Mar 07||
1. Recalibration of CO2 and O3 sensors. CO2 will
be rechecked in two days, as detected changes seem
2. 2005_only.xls file was reedited for CO2 data; comments added to meta_2005 sheet.
|09 Mar 07||Added solar related trends (total solar energy, sunshine, UVB and UVA)|
|06 Mar 07||Rechecked CO2 calibration.|
|05 Mar 07||Rechecked NOx and CO2 calibration. The old span gas bottle is empty; switched to a fresh 508ppm bottle. More calibrations will be needed during the coming days, so don't be afraid by spiky graphs.|
|Uploaded feb07.dat file into data archive. Please read header of file and look here!|
|03 Mar 07||Big maintenance by Envitec. NOx sensor repaired and restarted. Fault of CO2 sensor found and repaired. CO2 data from February will be rechecked and dubious ones discarded.|
|10 Feb 07||Maintenance work and recalibration. NO/NO2 sensor is shut down because of ozone generator and pump failure. So no NO/NO2 readings from today 15:30 on.|
|04 Feb 07||Added a link TRENDS in the live-data page and the data archive. This points to a page showing the linear trends over the last years (1998 -2006). Made some changes at 22:30 UTC: there was a scaling error.|
|03 Feb 07||Uploaded jan07.dat data file.|
|02 Feb 07||Finalized version 1.0 of CO2 paper.|
|22 Jan 07||Still struggling with CO2 sensor. Made zero check and recalibration; wait and see...|
|19 Jan 07||A few hours checking and recalibrating; sensor seems to be ok again... let's wait until next night.|
|18 Jan 07||Big Problem with the CO2 sensor: after midnight, readings plunge to ~120 ppm; impossible to recalibrate correctly (readings reamin below 300ppm). Problem could be caused by faulty Zero Gas valve, so that the Reference Zero made at night would be wrong (zero would be attributed to mixture of outside air and zero air), and as consequence all following readings would be too low.|
|17 Jan 07||
1. added SEARCH window (searching by www.opencrawl.de
; you may also do the search through this link by choosing
"Meteorologie (Luxemburg) in the left selection box); there
remain some teething problems
2. maintenance and recalibration work
|13 Jan 07||I am preparing ( with coauthors) a paper on our CO2 measurements since 2003 to 2005. This is the first paper of this kind published in Luxembourg. A prerelease version of part I is available here. Please feel free to make any comments!|
|06 Jan 07||Replaced main gas pump; the old one did not seem to be in bad shape, and I found no leaks.|
|02 Jan 07||Uploaded dec06.dat and dobson06.html files to data archive.|
|01 Jan 07||
1. The usual new year problem with the 7 days
graphs.. as the logger file does not contain the year, Gnuplot
gets confused the 1rst week of every new year. This will be
corrected the 7th January.
2. Main gas pump breakdown during 31Dec06; pump will be replaced the 6th Jan. Meanwhile take all CO2,O3,NO and NO2 readings as mean background values.